# ScaffoldGPT: A Scaffold-based Large Language Model for Drug Improvement

### v1

@@ Abstract
Drug optimization has become increasingly crucial in light of fast-mutating virus strains and drug-resistant cancer cells. Nevertheless, it remains challenging as it necessitates retaining the beneficial properties of the original drug while simultaneously enhancing desired attributes beyond its scope. In this work, we aim to tackle this challenge by introducing ScaffoldGPT, a novel Large Language Model (LLM) designed for drug optimization based on molecular scaffolds. Our work comprises three key components: (1) A three-stage drug optimization approach that integrates pretraining, finetuning, and decoding optimization. (2) A uniquely designed two-phase incremental training approach for pre-training the drug optimization LLM-based generator on molecule scaffold with enhanced performance. (3) A token-level decoding optimization strategy, Top-N, that enabling controlled, reward-guided generation using pretrained/finetuned LLMs. Finally, by conducting a comprehensive evaluation on COVID and cancer benchmarks, we demonstrate that ScaffoldGPT outperforms the competing baselines in drug optimization benchmarks, while excelling in preserving the original functional scaffold and enhancing desired properties.
Machine Learning, ICML
[…]
@@ 1 Introduction
The rise of rapidly mutating virus strains [14], exemplified by those related to SARS-CoV-2 [49], along with drug-resistant cancer cells [27], has heightened the urgency and interest in accelerating the development of effective treatments. However, traditional De Novo drug discovery processes are prohibitively expensive, often costing from hundreds of millions to billions of dollars [5], due to their extensive and resource-demanding nature. Despite considerable efforts, the success rate of drug discovery remains low, with many candidates failing early in the development process. This has led to an increased focus on drug repurposing, which leverages existing FDA-approved drugs for new therapeutic uses rather than creating new drugs from scratch. While drug repurposing has seen some success [34], its effectiveness is often constrained because drugs are typically developed with a high specificity for a particular disease.
@@
Drug improvement seeks to overcome the limitations inherent in both De Novo drug discovery and drug repurposing by enhancing an existing FDA-approved drug. Drug improvement is becoming an increasingly vital field, yet it remains relatively underexplored compared to drug discovery and repurposing efforts. DrugImprover [22] has started to clearly define the drug optimization problem by using Tanimoto similarity. Additionally, the DrugImprover framework contributes to the drug improvement domain in three key aspects: 1) a detailed workflow for drug optimization, 2) it introduced an Advantage-alignment Policy Optimization (APO) reinforcement learning (RL) algorithm to enhance the multi-objective generative model for drug improvement, and 3) includes an extensive dataset featuring 1 million ligands and their OEDOCK scores for five proteins related to human cancer cells and 24 high-affinity binding sites on the SARS-CoV-2 protein 3CLPro (PDB ID: 7BQY). The DrugImprover framework features a pretrained generative model that is subsequently refined using the APO reinforcement learning algorithm to ensure the molecules produced align with new objectives. Although DrugImprover has demonstrated encouraging outcomes, its effectiveness is limited due to its dependence on the less complex LSTM network architecture, which might lead to limited scalability and capacity, contextual understanding.
@@
On the other hand, LLMs have shown exceptional performance in diverse areas, from natural language understanding and generation exemplified by ChatGPT [48, 32], to text-to-video conversion as seen in SORA [25], and in programming and code generation through platforms like PG-TD [50] and Copilot [30]. Nevertheless, in the field of drug discovery, tools such as DrugGPT [21] and ChemGPT [9], despite making some initial strides, have not yet achieved performance on par with their counterparts in other domains. The field of drug discovery remains in expectation of breakthroughs similar to the ones achieved by LLMs in other domains.
@@
Several challenges have hindered the impact of large language models (LLMs) on drug design: Firstly, the molecules created need to satisfy multiple criteria such as solubility and synthesizability, and must also secure a high docking score against a specific target site. However, existing drug discovery LLMs generally only undergo pretraining with molecules and do not focus on enhancing multiple attributes simultaneously. Secondly, given that drugs sharing similar chemical structures tend to exhibit comparable biological or chemical effects [2], it is crucial to optimize the drug while preserving the beneficial chemical structure of the original molecule. Lastly, the current methodologies in drug discovery LLMs primarily focus on maximizing likelihood during the decoding phase, rather than customizing the optimization process to meet specific goals.
@@
REINVENT 4 [15, 16, 26] achieves the state of art performance in LLM-based drug optimization by using the Transformer model and conducting experiments with randomly selected ligand pairs that maintain constrained Tanimoto Similarity. While this method successfully produces ligands with high Tanimoto Similarity, it encounters difficulties in consistently improving the properties of the original drug and tends to simply maximize the likelihood of molecule generation.
@@
Our contributions. In this work, we propose ScaffoldGPT, a novel scaffold-based large language model with three-stage optimization process for drug improvement. This optimization process is designed to enhance existing drugs to rapidly evolving virus variants and cancer cells, overcoming the limitation of earlier drug improvement efforts. The contributions of ScaffoldGPT are summarized as follows:
@@
$\bullet$ A framework for drug improvement using LLMs that involves a three-step optimization process. The underlying motivation is that each stage is complementary to each other, enhances the performance of the preceding one. Furthermore, we have conducted ablation studies to illustrate the performance gains achieved at each step.
@@
$\bullet$ A scaffolds-based LLM with a novel two-phase incremental training specifically designed for drug optimization. The motivation for incremental training is to conduct local optimization before embarking on global optimization by breaking down the whole training corpus into knowledge pieces in an incremental order. We also conducted an ablation study in table 3 to demonstrate the effectiveness of the incremental training.
@@
$\bullet$ A novel token-level decoding optimization strategy, Top-N, utilizes pretrained LLMs to enable controlled, reward-guided generation that aligns with targeted objectives.
@@ 2 Related Work
2.1 Reinforcement Learning and RL finetuning
Reinforcement learning [41] has become a fundamental strategy in drug design [3, 13, 29, 31, 33, 39, 42, 45, 51, 52], focusing on optimizing rewards that aggregate predicted property scores from various pharmaceutical predictors. Traditional reinforcement learning methods in drug discovery often neglected molecular structure constraints, leading to significant structural changes and unsynthesizable compounds. Our research differs by refining existing drugs to enhance their attributes without redesigning them from scratch and using reinforcement learning to improve a pre-trained large language model generator, rather than starting anew. We also incorporate methods like Advantage-aligned Policy Optimization (APO) [22], which assigns rewards based on advantage preference over the original molecule, to fine-tune a Transformer model, ensuring it aligns with multiple pharmaceutical objectives while preserving molecular structure. This approach, which includes controllable decoding, refines the model beyond traditional reinforcement learning fine-tuning stages.
2.2 Planning with LLMs
Several studies have leveraged planning algorithms to enhance text outputs for a variety of NLP tasks. These include approaches like beam search optimization, machine translation improvements [37, 20, 4]. The PG-TD [50] method is tailored for code generation using a singular reward function, whereas ERP [23] introduces a novel concept by considering the certainty of each generated token along with an e-step forward entropy measurement to gauge potential outcomes. It has been shown that ERP effectively balances exploration and exploitation within molecular structures, leading to the discovery of high-reward molecules. Unlike previous studies that focus solely on planning with pre-trained large language models, our approach incorporates a novel decoding optimization as a critical final step in the algorithm. Moreover, our focus is on optimizing an existing drug rather than creating a new one from the ground up.
@@ 2.3 Drug Optimization
Recent efforts in drug optimization have primarily focused on a limited array of drug properties while often disregarding the docking score, an essential metric for evaluating structural compatibility with a target [52, 6]. DrugEx v3 [24] seeks to resolve this by using 3D molecular graphs that encapsulate more comprehensive data such as chemical valence rules. Nevertheless, this method’s complexity poses challenges in generating molecules that closely resemble the originals, thus leading to lower similarity. This reduced similarity could account for the diminished efficacy of graph-based methods, as deviations from the original molecular structures result in the loss of vital chemical properties. Conversely, our approach manages to preserve a significant level of similarity. DrugImprover [22] effectively begins to redefine the drug optimization challenge by employing reinforcement learning with a mix of multiple objectives as rewards. Still, it employs an LSTM in its generative model, which faces limitations in scalability, capacity, and understanding context. In contrast, our method utilizes a Transformer as the primary generative model, enhanced further with Advantage-aligned Policy Optimization (APO) for exploratory purposes and an optimized decoder for superior performance.
2.4 Large Language Model for Drug Discovery
@@
Large language models like MolGPT [1] and ChemGPT [9] have been utilized in molecule generation and drug discovery [1, 36, 46], using formats like SMILES [47] and SELFIES [18] to standardize molecular representation. These models have shown promise in drug design, outperforming traditional methods in some predictive capacities. However, as noted by  Murakumo et al. [28], using pre-trained large language models typically results in minor molecular modifications and serves primarily supportive roles in the design process. Our research adopts the SMILES format and combines the strengths of large language models with reinforcement learning. Unlike previous efforts that relied solely on LLMs, our approach uses RL fine-tuning to align with multiple pharmaceutical objectives and employs controlled decoding to guide LLMs in generating more effective molecular structures. Further developments include transformer-based models like REINVENT 4  [15, 16, 26], which mostly focuses on pretraining. While pretraining helps in generating molecules similar to those in the training dataset, it also inherently restricts the exploration scope due to biases in the training data. REINVENT4 uses the original molecule as input, generates molecules that are very similar to the original ones. This likely contributes to its relatively poor performance in other metrics, as the generated molecules exhibit minimal changes in chemical properties. In contrast, our proposed method, which uses SMILES and scaffolds as prompts, achieves a well-balanced trade-off between diversity and similarity in the generated and original molecules, potentially leading to improved performance.
@@ 3 Preliminaries
In the following sections, we detail MDP, LLM and drug discovery, complete with their mathematical notations, and integrate them within the framework of Markov decision processes.
Markov decision processes. Let us define a finite-horizon Markov decision process (MDP) [35] $\mathcal{M}_{0}=\langle\mathcal{S},\mathcal{A},T,\mathcal{P},R\rangle$. In this context, $\mathcal{S}$ represents a finite set of states, while $\mathcal{A}$ comprises a finite set of actions. The term $T$ denotes the planning horizon. The function $\mathcal{P}$, defined as $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}^{\prime}$, describes the deterministic transition dynamics that combine a state $s$ with an action $a$, with an episode concluding once the agent executes the termination action. Additionally, the reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scores exclusively to complete molecules, assigning a reward of 0 to partial molecules. The effectiveness of a policy can be evaluated using the $Q$-value function, denoted as $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and defined by the following equation: $Q^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{t=0}^{T}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right],$ where the expectation is based on the trajectory determined by the policy $\pi$. The associated value function is given by: $V^{\pi}(s):=\mathbb{E}_{a\sim\pi(\cdot|s)}\left[Q^{\pi}(s,a)\right]$.
@@
Limitation of previous works: DrugImprover, which utilizes LSTM networks, has limitations in scalability, capacity, and contextual understanding, especially when compared to versions that use Large Language Models. The current state of the art, REINVENT 4, employs the Transformer architecture; however, it mainly focuses on pretraining with constrained similarity, which restricts its capability to explore molecular spaces that might offer high rewards beyond its training set. In this work, we address these limitations by proposing ScaffoldGPT.

### v2

# ScaffoldGPT: A Scaffold-based GPT Model for Drug Optimization

@@ Abstract
Drug optimization has become increasingly crucial in light of fast-mutating virus strains and drug-resistant cancer cells. Nevertheless, it remains challenging as it necessitates retaining the beneficial properties of the original drug while simultaneously enhancing desired attributes beyond its scope. In this work, we aim to tackle this challenge by introducing ScaffoldGPT, a novel Generative Pretrained Transformer (GPT) designed for drug optimization based on molecular scaffolds. Our work comprises three key components: (1) A three-stage drug optimization approach that integrates pretraining, finetuning, and decoding optimization. (2) A uniquely designed two-phase incremental training approach for pre-training the drug optimization GPT on molecule scaffold with enhanced performance. (3) A token-level decoding optimization strategy, Top-N, that enabling controlled, reward-guided generation using pretrained/finetuned GPT. We demonstrate via a comprehensive evaluation on COVID and cancer benchmarks that ScaffoldGPT outperforms the competing baselines in drug optimization benchmarks, while excelling in preserving original functional scaffold and enhancing desired properties.
@@ 1 Introduction
The rise of rapidly mutating virus strains (Hadj Hassine, 2022), exemplified by those related to SARS-CoV-2 (Yuki et al., 2020), along with drug-resistant cancer cells (Mansoori et al., 2017), has heightened the urgency and interest in accelerating the development of effective treatments. However, traditional De Novo drug discovery processes are prohibitively expensive, often costing from hundreds of millions to billions of dollars (Dickson and Gagnon, 2009), due to their extensive and resource-demanding nature. Despite considerable efforts, the success rate of drug discovery remains low, with many candidates failing early in the development process. This has led to an increased focus on drug repurposing, which leverages existing FDA-approved drugs for new therapeutic uses rather than creating new drugs from scratch. While drug repurposing has seen some success (Pushpakom et al., 2019), its effectiveness is often constrained because drugs are typically developed with a high specificity for a particular disease.
@@
Drug optimization seeks to overcome the limitations inherent in both De Novo drug discovery and drug repurposing by enhancing an existing FDA-approved drug. Drug optimization is becoming an increasingly vital field, yet it remains relatively underexplored compared to drug discovery and repurposing efforts. DrugImprover (Liu et al., 2023a) has started to clearly define the drug optimization problem by using Tanimoto similarity. Additionally, the DrugImprover framework contributes to the drug optimization domain in three key aspects: a detailed workflow for drug optimization, an Advantage-alignment Policy Optimization (APO) reinforcement learning (RL) algorithm to enhance the multi-objective generative model for drug optimization, and an extensive dataset featuring 1 million ligands and their OEDOCK scores for five proteins related to human cancer cells and 24 high-affinity binding sites on the SARS-CoV-2 protein 3CLPro (PDB ID: 7BQY). The DrugImprover framework features a pretrained generative model that is subsequently refined using the APO reinforcement learning algorithm to ensure the molecules produced align with new objectives. Although DrugImprover has demonstrated encouraging outcomes, its effectiveness is limited due to its dependence on the less complex LSTM network architecture, which might lead to limited scalability and capacity, contextual understanding.
@@
On the other hand, generative pretrained language models have shown exceptional performance in diverse areas, from natural language understanding and generation exemplified by ChatGPT (Wu et al., 2023, Ouyang et al., 2022), to text-to-video conversion as seen in SORA (Liu et al., 2024c), and in programming and code generation through platforms like PG-TD (Zhang et al., 2023a) and Copilot (Nguyen and Nadi, 2022). Nevertheless, in the field of drug discovery, tools such as DrugGPT (Li et al., 2023), ChatDrug (Liu et al., 2024a) and ChemGPT (Frey et al., 2023), despite making some initial strides, have not yet achieved performance on par with their counterparts in other domains. The field of drug discovery remains in expectation of breakthroughs similar to the ones achieved by LMs in other domains.
@@
Several challenges have hindered the impact of Language Models (LMs) on drug design: Firstly, the molecules created need to satisfy multiple criteria such as solubility and synthesizability, and must also secure a high docking score against a specific target site. However, existing drug discovery LMs generally only undergo pretraining with molecules and do not focus on enhancing multiple attributes simultaneously. Secondly, given that drugs sharing similar chemical structures tend to exhibit comparable biological or chemical effects (Bender and Glen, 2004), it is crucial to optimize the drug while preserving the beneficial chemical structure of the original molecule. Lastly, the current methodologies in drug discovery LMs primarily focus on maximizing likelihood during the decoding phase, rather than customizing the optimization process to meet specific goals.
@@
REINVENT 4 (He et al., 2021; 2022, Loeffler et al., 2024) achieves the state of art performance in LM-based drug optimization by using the Transformer model and conducting experiments with randomly selected ligand pairs that maintain constrained Tanimoto Similarity. While this method successfully produces ligands with high Tanimoto Similarity, it encounters difficulties in consistently improving the properties of the original drug and tends to simply maximize the likelihood of molecule generation.
@@
In this work, we propose ScaffoldGPT, a novel scaffold-based GPT with three-stage optimization process for drug optimization. This optimization process is designed to enhance existing drugs to rapidly evolving virus variants and cancer cells, overcoming the limitation of earlier drug optimization efforts. The contributions of ScaffoldGPT are summarized as follows:
@@
$\bullet$ A framework for drug optimization using GPT that involves a three-step optimization process. The underlying motivation is that each stage is complementary to each other, enhances the performance of the preceding one. Furthermore, we have conducted ablation studies to illustrate the performance gains achieved at each step.
@@
$\bullet$ A scaffolds-based GPT with a novel two-phase incremental training specifically designed for drug optimization. The motivation for incremental training is to conduct local optimization before embarking on global optimization by breaking down the whole training corpus into knowledge pieces in an incremental order. We also conducted an ablation study in table 3 to demonstrate the effectiveness of the incremental training.
@@
$\bullet$ A novel token-level decoding optimization strategy, Top-N, utilizes pretrained GPT to enable controlled, reward-guided generation that aligns with targeted objectives.
@@ 2 Related Work
2.1 Reinforcement Learning and Finetuning
Reinforcement learning (Tan et al., 2022a) has become a fundamental strategy in drug design (Born et al., 2021, Guimaraes et al., 2017, Neil et al., 2018, Olivecrona et al., 2017, Popova et al., 2018, Ståhl et al., 2019, Tan et al., 2022b, Wang et al., 2022a, Zhang et al., 2023b, Zhou et al., 2019), focusing on optimizing rewards that aggregate predicted property scores from various pharmaceutical predictors. Traditional reinforcement learning methods in drug discovery often neglected molecular structure constraints, leading to significant structural changes and unsynthesizable compounds. Our research differs by refining existing drugs to enhance their attributes without redesigning them from scratch and using reinforcement learning to improve a pre-trained language model generator, rather than starting a new. We also incorporate methods like Advantage-aligned Policy Optimization (APO) (Liu et al., 2023a), which assigns rewards based on advantage preference over the original molecule, to fine-tune a Transformer model, ensuring it aligns with multiple pharmaceutical objectives while preserving molecular structure. This approach, which includes controllable decoding, refines the model beyond traditional reinforcement learning fine-tuning stages.
2.2 Planning with GPT Models
Several studies have leveraged planning algorithms to enhance text outputs for a variety of NLP tasks. These include approaches like beam search optimization, machine translation improvements (Scialom et al., 2021, Leblond et al., 2021, Chaffin et al., 2021). The PG-TD (Zhang et al., 2023a) method is tailored for code generation using a singular reward function, whereas ERP (Liu et al., 2024b) introduces a novel concept by considering the certainty of each generated token along with an e-step forward entropy measurement to gauge potential outcomes. It has been shown that ERP effectively balances exploration and exploitation within molecular structures, leading to the discovery of high-reward molecules. Unlike previous studies that focus solely on planning with pre-trained language models, our approach incorporates a novel decoding optimization as a critical final step in the algorithm. Moreover, our focus is on optimizing an existing drug rather than creating a De Novo one from the ground up.
@@ 2.3 Drug Optimization
Recent drug optimization efforts have focused primarily on a limited array of drug properties while often disregarding the docking score, an essential metric for evaluating structural compatibility with a target (Zhou et al., 2019, Erikawa et al., 2021). DrugEx v3 (Liu et al., 2023b) seeks to resolve this deficiency by using 3D molecular graphs that encapsulate more comprehensive data such as chemical valence rules. Nevertheless, this method’s complexity poses challenges in generating molecules that closely resemble the originals. The resulting lower similarity could account for the diminished efficacy of graph-based methods, as deviations from the original molecular structures result in the loss of vital chemical properties. Conversely, our approach manages to preserve a decent level of similarity. Diffusion-based efforts (Alakhdar et al., 2024, Morehead and Cheng, 2024) either focus on De Novo drug discovery while neglecting the essential similarity to the original molecule, or overlook the docking score towards a binding target. Molsearch (Sun et al., 2022) is a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization, and MIMOSA (Fu et al., 2021) is a GNN sampling-based method leveraging graph-based molecular optimization. DrugImprover (Liu et al., 2023a) effectively begins to redefine the drug optimization challenge by employing reinforcement learning with a mix of multiple objectives as rewards. Still, it employs an LSTM in its generative model, which faces limitations in scalability, capacity, and understanding context. In contrast, our method applies a Transformer as the primary generative model, enhanced further with Advantage-aligned Policy Optimization (APO) for exploratory purposes and an optimized decoder for superior performance.
2.4 Language Models for Drug Discovery
@@
Large language models like MolGPT (Bagal et al., 2021) and ChemGPT (Frey et al., 2023) have been utilized in molecule generation and drug discovery (Bagal et al., 2021, Rothchild et al., 2021, Wang et al., 2022b), using formats like SMILES (Weininger, 1988) and SELFIES (Krenn et al., 2020) to standardize molecular representation. These models have shown promise in drug design, outperforming traditional methods in some predictive capacities. However, as noted by  Murakumo et al. (2023), using pre-trained language models typically results in minor molecular modifications and serves primarily supportive roles in the design process. We adopt the SMILES format and combine the strengths of GPTs with reinforcement learning. Unlike previous efforts that relied solely on LLMs, our approach uses RL fine-tuning to align with multiple pharmaceutical objectives and employs controlled decoding to guide the GPT model in generating more effective molecular structures.
Further developments include transformer-based models like REINVENT 4  (He et al., 2021; 2022, Loeffler et al., 2024), which mostly focuses on pretraining. While pretraining helps in generating molecules similar to those in the training dataset, it also inherently restricts the exploration scope due to biases in the training data. REINVENT4 uses the original molecule as input, generates molecules that are very similar to the original ones. This likely contributes to its relatively poor performance in other metrics, as the generated molecules exhibit minimal changes in chemical properties. In contrast, our proposed method, which uses SMILES and scaffolds as prompts, achieves a well-balanced trade-off between diversity and similarity in the generated and original molecules, potentially leading to improved performance.
@@ 3 Preliminaries
In the following sections, we detail MDP, Language Model (LM) and drug discovery, complete with their mathematical notations, and integrate them within the framework of Markov decision processes.
Markov decision processes. Let us define a finite-horizon Markov decision process (MDP) (Puterman, 2014) $\mathcal{M}_{0}=\langle\mathcal{S},\mathcal{A},T,\mathcal{P},R\rangle$. In this context, $\mathcal{S}$ represents a finite set of states, while $\mathcal{A}$ comprises a finite set of actions. The term $T$ denotes the planning horizon. The function $\mathcal{P}$, defined as $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}^{\prime}$, describes the deterministic transition dynamics that combine a state $s$ with an action $a$, with an episode concluding once the agent executes the termination action. Additionally, the reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scores exclusively to complete molecules, assigning a reward of 0 to partial molecules. The effectiveness of a policy can be evaluated using the $Q$-value function, denoted as $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and defined by the following equation: $Q^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{t=0}^{T}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right],$ where the expectation is based on the trajectory determined by the policy $\pi$. The associated value function is given by: $V^{\pi}(s):=\mathbb{E}_{a\sim\pi(\cdot|s)}\left[Q^{\pi}(s,a)\right]$.
@@
Limitation of previous works: DrugImprover, which utilizes LSTM networks, has limitations in scalability, capacity, and contextual understanding, especially when compared to versions that use GPT Models. The current state of the art, REINVENT 4, employs the Transformer architecture; however, it mainly focuses on pretraining with constrained similarity, which restricts its capability to explore molecular spaces that might offer high rewards beyond its training set. In this work, we address these limitations by proposing ScaffoldGPT.
@@ 4 The ScaffoldGPT Algorithm
1: GPT-based generator policy $\pi_{{\theta}}$; critics $\mathbf{{C}}$.
2:Initialize $\pi_{{\theta}}$ with GPT2-like Transformer with random weight $\theta$.
3$\triangleright$ /* Stage 1: Pretrain GPT-based generator */
4:Build the training corpus $C_{\text{Phase 1}}$ (4), $C_{\text{Phase 2}}$ (5).
5:Pre-train BPE tokenizer and $\pi_{{\theta}}$ on $C_{\text{Phase 1}}$ via CLM objective (3).
6:Pre-train $\pi_{{\theta}}$ on $C_{\text{Phase 2}}$.
7$\triangleright$ /* Stage 2: APO fine-tuning */
8:for $n=1,\ldots,N$ do
9:     $s_{0}\sim\rho_{0}$, where $\rho_{0}\in\Delta\left({\mathcal{B}}\right)$.
10:     Generate $Y_{1:T}=\left({y_{t},\ldots,y_{T}}\right)\sim\pi_{{\theta}}\left({\cdot|S}\right)$.
11:     Compute advantage preference $R^{\text{AP}}$ by (9)(10).
12:     Update generator $\theta$ via policy gradient by (11).
13$\triangleright$ /* Stage 3: Token-level Decoding Optimization */
14:Optimize the generation of $\pi_{{\theta}}$ via Top-N (12) decoding strategy.
Stage 1. Pretrain a GPT generator. Let us note the GPT generator policy as $\pi_{\theta}$, which computes the probability $p$ of the occurrence of the $t^{th}$ token in a target molecule $Y$. It takes into account all preceding tokens $\mathbf{y}_{<t}=\left[y_{1},...,y_{t-1}\right]$ in the target, as well as the scaffold compound $S$, which is noted as $\pi_{\theta}\left({y_{t}{\,|\,}\mathbf{y}_{<t},S}\right)=p\left({y_{t}|\mathbf{y}_{<t},S}\right).$ The parameters $\theta$ of the generator policy $\pi_{\theta}$ are trained using the training corpus set through the minimization of the negative log-likelihood (NLL) for the complete SMILES strings across the entire set. This process is described as follows:
$NLL=-\log{P}\left({Y|S}\right)=-\sum_{t=1}^{T}\log{P}\left({y_{t}|y_{t-1},...,y_{1},S}\right)=-\sum_{t=1}^{T}\log\pi_{\theta}\left({y_{t}|y_{1:t-1},S}\right),$ (3)
In this project, we employ pre-training to harness large quantities of unlabeled text to construct a basic foundation model of language understanding. This foundation model can subsequently be fine-tuned and tailored to meet various specialized goals. In this work, we propose a novel framework for pre-training a Transformer and linking scaffolds with complete molecules, based on a SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988) string representation of the molecule. Here, we define $\mathcal{S}$ as the scaffold space and $\mathcal{Y}$ as the target molecule space. With $\mathcal{P}=\left\{\left({s,y}\right)|s,y\in\mathcal{S}\times\mathcal{Y}\right\}$, we denote the set of scaffold and molecular pairs from $\mathcal{S}$ and $\mathcal{Y}$. In this notation, $s$ represents the scaffold of the target molecule, and $y$ is the corresponding target molecule. We initially pre-trained our tokenizer using the Byte Pair Encoding (BPE) method (Gage, 1994, Sennrich et al., 2015). Building on the pre-trained BPE tokenizer, we propose a two-phase incremental training approach, as illustrated in Fig. 1, to notably enhance the model’s ability to improve the validity of inferring the target molecule from its scaffold.
Incremental training.
The rationale for incremental training is to conduct local optimization before embarking on global optimization. Therefore, we divide our training into two phases. In the first phase, we focus on training a GPT exclusively for molecules using Causal Language Modeling (CLM). CLM utilizes an autoregressive method where the model is trained to predict the next token in a sequence by considering only the tokens that precede it. The phase 1 corpus is designed as follows:
$C_{\text{Phase 1}}=\left\{[BOS],\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},\left\langle L\right\rangle\text{is the token for ligand.}$. (4)
In the second phase, building upon the success of the GPT model developed in phase 1, which demonstrated high accuracy in molecular generation, we advance the training by focusing on pairs of scaffolds and molecules using CLM. The phase 2 corpus is as follows:
$C_{\text{Phase 2}}=\left\{[BOS],\left\langle S\right\rangle,\underbrace{s_{1},\cdots,s_{T}}_{\text{source scaffold S}},\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},\left\langle S\right\rangle\text{: scaffold.}$. (5)
Consequently, the model can generate the appropriate molecule when given a scaffold. However, since a single scaffold may correspond to multiple molecules, we further refine the GPT-based generator policy, $\pi_{\theta}$, to target specific outcomes by applying reinforcement learning finetuning in the next stage.
Stage 2. RL finetuning. Fine-tuning a generative model is crucial for producing outcomes that meet specific objectives. In this study, we use the Advantage-alignment Policy Optimization (APO) (Liu et al., 2023a) algorithm to fine-tune the pretrained GPT-based generator policy $\pi_{\theta}$. This approach steers the model from a given scaffold towards the targeted molecule, simultaneously improving multiple properties.
In this work, we adopt the define of reward function from Liu et al. (2024b), and regarded each pharmaceutical property as a critic $C$ and got an ensemble critics $\mathbf{C}$ as follows:
${R^{\text{norm}}\left({Y}\right)}:={R^{\text{norm}}\left({Y|S}\right)}=\lambda\cdot\text{Norm}\left({{C}^{\text{Tanimoto}}\left({S,Y}\right)}\right)+\sum_{i=0}^{|\mathbf{C}|-1}\lambda\cdot\text{Norm}\left({{C}_{i}{\left({Y}\right)}}\right)$ (6)
where Norm is employed to standardize diverse attributes to a consistent scale. In this instance, Norm refers to the process of min-max normalization, which is used to adjust the attributes so they fit within the $\left[0,1\right]$ range. We defer the details of critics to §5.1.
In this work, we use BON (Gao et al., 2023) (Best of N) search to estimate the reward for a prompt (partial molecule). BON can be formulated as the following:
where $\text{where }Y_{j}=$ $\left[\mathbf{y}_{<i},y_{i},\cdots,y_{T}\right]_{j},\text{and }y_{i}\sim\textsc{Top-PK}\left({\mathbf{y}_{<i},p,k}\right)|_{S}.$
where Top-PK (Liu et al., 2024b) is defined as follows:
[…]
$j=\min\left\{\mathop{\rm arg\,min}_{j^{\prime}}\sum_{i=1}^{j^{\prime}}\pi_{\theta}\left({y_{i}|S,\mathbf{y}_{<i}}\right)\geq p,k\right\},\text{and }\pi_{\theta}\left({y_{g}|S,\mathbf{y}_{<i}}\right)>\pi_{\theta}\left({y_{h}|S,\mathbf{y}_{<i}}\right),\text{if }g<h,$
APO makes policy gradient based on the advantage preference (Liu et al., 2023a), which is defined as
and perform APO policy gradient with follows:
Stage 3. Token-level Controllable decoding generation.
Ultimately, the current GPT-based decoder focuses mainly on maximizing likelihood, neglecting specific metrics of interest. This approach limits its effectiveness in optimizing objectives that diverge from those in its training set, particularly in generating desired molecules. In this study, we introduce controllable decoding after fine-tuning with APO. We present a new approach, Top-N, to direct the generation process towards enhancements in the optimization objective, as detailed below: